OpenStudy (anonymous):

Factor completely. 40x^3-5y^3 Cn you please show work so I can understand how to do it :)

OpenStudy (anonymous):

40x^3-5y^3 = 5( 8x^3 - y^3 ) = 5 ( (2x)^3 - y^3 ) Now we use this identity: http://www.mathsisfun.com/algebra/polynomials-difference-two-cubes.html a^3 - b^3 = (a-b)(a^2+ab+b^2) In our case: a=2x b=y 5( (2x)^3 - y^3 ) = 5 (2x-y)(4x^2+2xy+y^2) More about factoring, and special identities: http://en.wikipedia.org/wiki/Difference_between_two_cubes#Sum.2Fdifference_of_two_cubes